# From paintbrush to pixel: A review of deep neural networks in AI-generated art

### v2

@@ 5 Ethical concerns
The ability of generative models to create images that are indistinguishable from real human-made images comes with several ethical concerns. These so-called ‘Deepfakes’ have caused harm in many ways Mirsky and Lee (2021); Hancock and Bailenson (2021); Diakopoulos and Johnson (2021). As a result, OpenAI and Google did not release their powerful text-to-image models fully open source. Users can only use filtered versions of their models through a web interface. These filtered versions have been trained on datasets that exclude potentially harmful images depicting violence, nudity or recognizable people (e.g., celebrities, political figures).
@@
In the context of art, these Deepfakes raise numerous other concerns Cejudo (2021). These models are often praised for their ability to replicate specific painters’ style or art movements. The flipside of the coin is that this blurs the borders of copyright and raises questions regarding authorship of AI mediated art. These models allow anyone to generate images in the style of an artist, without proper copyright protection. Shan et al. (2023) attempted to protect artists and their work by creating GLAZE, a model that allows artists to add an invisible watermark to their work that would distort a text-to-image model in learning their artistic style. As such, artists can protect their future work from copyright strikes. It is important to note that DALL-E 3 addresses concerns in the creative community by refusing requests to generate images in the style of living artists. It also maintains an updated blocklist of these artists’ names. For a more thorough review on copyright infringement in relation to AI-generated art, we refer the reader to Jiang et al. (2023).
@@ 6 Conclusions
Deep learning and its image processing applications are now at a totally different stage than they were a few short years ago. In the beginning of the last decade, it was groundbreaking that deep neural networks could classify natural images. Today, these models are capable of generating highly realistic and complex images based on simple text prompts. This allows individuals without programming knowledge to employ these powerful models. It is important to remember that the use of these models should be guided by ethical and responsible considerations. These tools can assist artists in expressing their creativity and may shape the future of art. It is undeniable that AI is affecting authorship and copyright rules regarding art, which comes with many ethical considerations. As ChatGPT stated, “Some people believe that AI has the potential to revolutionize the way we create and think about art, while others are more skeptical and believe that true creativity and artistic expression can only come from human beings. Ultimately, the role of AI in the arts will depend on how it is used and the goals and values of the people who are using it.”
Acknowledgments

### v3

@@ 5 Ethical concerns
The ability of generative models to create images that are indistinguishable from real human-made images comes with several ethical concerns. These so-called ‘Deepfakes’ have caused harm in many ways, including political misinformation, fraud through impersonation, and a broader loss of trust in digital media [110, 111, 112]. As a result, OpenAI and Google did not release their powerful text-to-image models fully open source. Users can only use filtered versions of their models through a web interface. These filtered versions have been trained on datasets that exclude potentially harmful images depicting violence, nudity or recognizable people (e.g., celebrities, political figures).
@@
In the context of art, these Deepfakes raise numerous other concerns [113]. These models are often praised for their ability to replicate specific painters’ style or art movements. The flipside of the coin is that this blurs the borders of copyright and raises questions regarding authorship of AI mediated art. These models allow anyone to generate images in the style of an artist, without proper copyright protection. [114] attempted to protect artists and their work by creating GLAZE, a model that allows artists to add an invisible watermark to their work that would distort a text-to-image model in learning their artistic style. As such, artists can protect their future work from copyright strikes. It is important to note that DALL-E 3 addresses concerns in the creative community by refusing requests to generate images in the style of living artists. It also maintains an updated blocklist of these artists’ names. For a more thorough review on copyright infringement in relation to AI-generated art, we refer the reader to [115].
@@ 6 Conclusions
Deep learning and its image processing applications are now at a totally different stage than they were a few short years ago. In the beginning of the last decade, it was groundbreaking that deep neural networks could classify natural images. Today, these models are capable of generating highly realistic and complex images based on simple text prompts. This allows individuals without programming knowledge to employ these powerful models. It is important to remember that the use of these models should be guided by ethical and responsible considerations. AI-generated art also invites reflection on broader issues, such as potential applications in art therapy and how creativity and authorship are understood. These tools are increasingly being explored by artists and they may influence artistic practice and discussions about the future of art. It is undeniable that AI is affecting authorship and copyright rules regarding art, which comes with many ethical considerations. As ChatGPT stated, “Some people believe that AI has the potential to revolutionize the way we create and think about art, while others are more skeptical and believe that true creativity and artistic expression can only come from human beings. Ultimately, the role of AI in the arts will depend on how it is used and the goals and values of the people who are using it.”
Acknowledgements
@@
Author contributions statement
A-S. Maerten and D.Soydaner contributed equally to this work. All authors reviewed the manuscript.